OpenStudy (anonymous):

Verify that y=tan(x+c) is a one parameter family of solutions of the differential equation y'=1+y^2

OpenStudy (anonymous):

y=tan(x+c) y'=sec^2(x+c)

OpenStudy (anonymous):

Answer: No real solutions.

OpenStudy (irishboy123):

keep going @Mateaus

OpenStudy (anonymous):

y'=sec^2(x+c) y'=tan^2(x+c)+1 that's how far I can simplify it but I am unsure if there is a possible way to get y'=1+y^2

OpenStudy (irishboy123):

?? you have done it, you were given : y'=1+y^2 with: y=tan(x+c) you showed that : y'=sec^2(x+c) [= y^2 +1]

OpenStudy (anonymous):

oh! derp... I see. Did the same mistake for another problem lol

OpenStudy (anonymous):

Thanks!